Circular No. 8338

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(22) KALLIOPE I (LINUS) J. Berthier, P. Descamps, and D. Hestroffer, Institut de Mécanique

Céleste et de Calcul des Éphémérides, Observatoire de Paris (OP); J. Lecacheux, Laboratoire d’Etudes Spatiales et d’Instrumentation en Astrophysique, OP; and F. Marchis, University of California at Berkeley, report on the probable ground track of two stellar occultations of Hipparcos/Tycho stars by Linus, the satellite of (22) Kalliope (cf. IAUC 7703, 8177). On May 14.5083 and June 5.8812 UTC, the stars TYC 6816-00087 (α = 17^{h}17^{m}.0, δ = −25^{o}08′, equinox 2000.0) and 6814-00458 (α = 16^{h}57^{m}.3, δ = −26^{o}04′) will be occulted for 15.4 s by (22) Kalliope with drops of 0.7 and 0.9 magni-

tude, respectively (Goffin 2003, ftp://ftp.ster.kuleuven.ac.be/dist/
vvs/asteroids; Preston 2004, http://www.asteroidoccultation.com/).

According to the orbit of Linus, as computed by Marchis et al. (2003, Icarus 165, 112) and improved with observations gathered during the past three years with several adaptive-optics systems (Lick 3-m, Palomar 5-m, VLT 8-m, and Keck 10-m telescopes), in addition to observations published by Margot and Brown (2003, Science 300, 1939), the two stellar occultations will occur when Linus is located, relative to the primary, at Δα = −0′′.14 and Δδ = +0′′.53 on May 14.5083, and at Δα = −0′′.47 and Δδ = −0′′.24 on June 5.8812. The occultations will be visible along a 40-km-wide shadow located, respectively, ∼ 8^{o} north and ∼ 5^{o} south of the primary occultation track, respectively, which itself crosses southern Australia and the south Pacific on May 14 and the northern Mediterranean Sea countries and Russia on June 5. The occultations by Linus are expected to last < 3.2 s. Shadow path, details, and updates can be obtained at website URL

http://www.imcce.fr/observateur/target/targobs.php?target=
Kalliope.

SUPERNOVAE 2004bn AND 2004bo T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 340–720 nm) of SN 2004bn (cf. IAUC 8335), obtained by P. Berlind on May 11.18 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova; adopting the NASA/IPAC Extragalactic Database recession velocity of 6533 km/s for the host galaxy, the expansion velocity, derived from the minimum of the Hβ line, is 5400 km/s. A spectrum of SN 2004bo (ibid.), obtained by Berlind on May 11.23, shows it to be a type-Ia supernova a few weeks past maximum.

2004 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green